OpenStudy (anonymous):

Find the cost of operating a 980-watt stereo system for 4 1/2 hours if he cost of electricity is 0.058 per kilowatt hour

OpenStudy (anonymous):

well, multiply 980 Watts by 4.5 hours to get ...?

OpenStudy (kropot72):

Convert 980 watts to kilowatts as follows: \[\frac{980\ watts}{1000}=0.98\ kilowatts\] Next multiply 0.98 kilowatts by 4.5 to get the quantity of kilowatt-hours. \[0.98\ kW \times 4.5\ hours=(0.98\times 4.5)\ kilowatt-hours\] Now multiply the number of kilowatt-hours by $0.058 to get the cost of operating the stereo: \[Cost = 0.98\times 4.5\times 0.058\ dollars\]